# Momentum-space Metric Tensor from Nonadiabatic Evolution of Bloch Electrons

### v1

@@ Abstract
We reveal a fundamental geometric structure of momentum space arising from the nonadiabatic evolution of Bloch electrons. By extending semiclassical wave packet theory to incorporate nonadiabatic effects, we introduce a momentum-space metric tensor—the nonadiabatic metric. This metric gives rise to two velocity corrections, dubbed geometric and geodesic velocities, providing a unified and intuitive framework for understanding nonlinear and nonadiabatic transport phenomena beyond Berry phase effects. Furthermore, we show that the nonadiabatic metric endows momentum space with a curved geometry, recasting wave packet dynamics as forced geodesic motion. In this picture, the metric defines distances, the Berry connection acts as a gauge potential, band dispersion serves as a scalar potential, and the toroidal topology of the Brillouin zone imposes periodic boundary conditions. When the nonadiabatic metric is constant, it reduces to an effective mass, allowing electrons to behave as massive particles in flat bands. In a flat Chern band with harmonic attractive interactions, the two-body wave functions mirror the Landau-level wave functions on a torus.
The adiabatic approximation, pioneered by Born and Oppenheimer, is a cornerstone of condensed matter physics. The emergence of Berry phase from adiabatic evolution [1] has profoundly shaped the study of topological phases of matter [2]. Adiabatic wave packet dynamics linked the Berry curvature to anomalous velocity in solids [3, 4, 5, 6], provided a unified geometric framework to understand topological phenomena [2]. This framework has revolutionized our understanding of how topological quantum materials respond to electric, magnetic, thermal, and lattice deformation stimuli [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
@@
Recent interest has surged in the Bloch-electron dynamics beyond Berry phase effects manifested in, for example, nonlinear responses [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. These phenomena are thought to stem from the quantum metric [23, 24, 25]—the real counterpart of the Berry curvature in the quantum geometric tensor [37]. The quantum metric is connected to rich physical effects, such as massive particle dynamics in flat bands [38, 39, 40, 41, 42, 43, 44, 45], exciton spectra [46, 47], fractional Chern insulators [48, 49, 50, 51], Wannier function spread [52], orbital magnetic susceptibility [53, 54], among others [23, 24, 25]. Despite its broad relevance, an intuitive framework for understanding those phenomena is still lacking.
In this Letter, we deliver a semiclassical perspective of the above phenomena by studying the nonadiabatic effects—the leading-order corrections to the Berry phase effects. While nonadiabatic effects have been explored in quantum systems driven by external classical parameters [55, 56, 57], their significance in the context of Bloch electrons remains elusive. Unlike the quantum-classical hybrid systems, Bloch electrons form a genuine quantum system, free from external classical degrees of freedom. Moreover, Bloch wavefunctions are parameterized by crystal momentum within the Brillouin zone—a torus characterized by nontrivial topology. This topological structure not only gives rise to the Berry phase in one dimension, known as the Zak phase [58], but also imposes periodic boundary conditions on electron dynamics within the parameter space, as we will demonstrate later.
Here, we extend semiclassical wave packet theory to incorporate nonadiabatic effects by including interband contributions (see Fig. 1). We begin by deriving the effective Lagrangian for the wave packet, revealing that nonadiabatic corrections emerge as a metric tensor in momentum space, which we term the nonadiabatic metric. We then discuss the relationship between this metric and the quantum metric and elucidate its profound implications for wave packet velocity, which provides a unified understanding of nonlinear and nonadiabatic transport. Next, we demonstrate that the nonadiabatic metric endows the Brillouin zone with the structure of a curved manifold, causing electron dynamics to resemble forced geodesic motion on this curved space. We further requantize the electron’s dynamics in momentum space, where the toroidal geometry of the Brillouin zone imposes periodic boundary conditions. When applied to flat Chern bands, this approach reveals the reemergence of massive particle dynamics and establishes a direct connection between bound-state wave functions and quantum Hall effects on a torus.
@@
Our results indicate that, in the dynamics of the Bloch electrons, the nonadiabatic metric is more fundamental than the quantum metric: it emerges directly from the Schrödinger equation as the leading correction beyond adiabatic evolution. It governs the dynamical behavior of Bloch electrons, providing a unified description of various dynamical phenomena. In contrast, the quantum geometric tensor is defined in the adiabatic limit, where it describes the geometric properties of the static Hilbert space [37].
Nonadiabatic Wave Packet Theory— We use the following wave packet ansatz [59]:
[…]
@@
with $|u_{n}(\boldsymbol{q})\rangle$ the periodic part of the Bloch state for band $n$, $c_{n}$ the coefficient, and $\sum_{n}|c_{n}(\boldsymbol{q},t)|^{2}=1$. The wave packet center is at $\bm{x}_{c}=\langle\Psi|\hat{\bm{x}}|\Psi\rangle$.
@@
Here we consider an isolated band labeled by index 0. We focus on the abelian case with $|c_{0}|\approx 1$ and $c_{n}\sim 0$ for $n\neq 0$. In contrast to the adiabatic wave packet dynamics developed by Sundaram and Niu [5], which constrains the wave packet on a single energy band, here we allow for interband contributions with $c_{n}\neq 0$ (see Fig. 1). These interband components are expressed as an asymptotic series in the frequency of external perturbations—a notable distinction from the higher-order theory in Ref. [27], where the wave packet is expanded as an asymptotic series in the strength of external fields.
@@
The dynamics of the wave packet is obtained by minimizing the Dirac-Frenkel action [60, 61]
[…]
where $\bm{x}_{c}$, $\bm{q}_{c}$, $c_{n}$, and $c_{n}^{*}$ are dynamical variables. By integrating out the dynamics of $c_{n}$ [59], we obtain our central result: the nonadiabatic effective Lagrangian (with $\hbar=1$)
$\mathcal{L}_{\rm eff}$ $=\mathcal{L}_{\rm eff,0}+\frac{1}{2}G_{ij}\dot{q}_{c}^{i}\dot{q}_{c}^{j}$ (4)
$\mathcal{L}_{\rm eff,0}$ $=-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{0}(\bm{q}_{c})-V(\bm{x}_{c})+\dot{q}_{c}^{i}A_{i}$
where $\mathcal{L}_{\rm eff,0}$ is the adiabatic effective Lagrangian with ${A}_{i}=\langle u_{0}|i\partial_{{q}^{i}}|u_{0}\rangle|_{\bm{q}_{c}}$ the Berry connection, $E_{n}$ the eigenenergy of the $n$-th band, and $V(\bm{x}_{c})$ the slowly varying external potential. The Einstein summation convention is implied throughout the manuscript unless otherwise noted. From now on, we drop the subscript $c$, using $(\bm{x},\bm{q})$ for $(\bm{x}_{c},\bm{q}_{c})$.
@@
Our key contribution is the identification of a new term in the effective Lagrangian, characterized by the coefficients $G_{ij}$. $G_{ij}$ acts as a metric tensor in momentum space, which is termed the nonadiabatic metric, as this metric originates from nonadiabatic corrections to wave packet dynamics: $c_{n}\propto\dot{q}^{i}$ for $n\neq 0$, which reflect a finite occupation of other instantaneous eigenstates. The matrix element reads
$G_{ij}=$ $2\text{Re}\sum_{m\neq 0}\frac{A_{i,[0m]}A_{j,[m0]}}{\omega_{[m0]}}$ (5)
@@
where $A_{i,[mn]}=\langle u_{m}|i\partial_{q^{i}}|u_{n}\rangle$, $\omega_{[mn]}=E_{m}-E_{n}$. The metric is a gauge-invariant symmetric tensor [62]. The nonadiabatic metric is related to but different from the quantum metric $g_{ij}=\text{Re}\mathcal{Q}_{ij}$ that is the real part of the quantum geometric tensor (or the Fubini-Study metric) $\mathcal{Q}_{ij}=\sum_{m\neq n}{A_{i,[nm]}A_{j,[mn]}}$. In a two-band model with an energy gap $\Delta(\bm{q})$, $G_{ij}$ of the lower band is proportional to $g_{ij}$: $G_{ij}=\frac{2}{\Delta}g_{ij}$.
@@
Just as adiabatic theory unifies Berry phase effects through the term $\dot{q}^{i}A_{i}$ [2], the nonadiabatic theory introduces the new term, $\frac{1}{2}G_{ij}\dot{q}^{i}\dot{q}^{j}$, providing a comprehensive framework for capturing a wide array of phenomena that extend beyond the Berry phase effects. As we show below, it is the nonadiabatic metric, rather than the quantum metric, that fundamentally governs a broad class of nonlinear and nonadiabatic transport phenomena.
At a deeper level, the nonadiabatic metric uncovers the intrinsically curved geometric structure of the Brillouin zone ($\bm{q}$-space) and highlights its topological character as a torus, as detailed later. Notably, this $\bm{q}$-space representation enables the reemergence of massive particle dynamics in flat bands with corrections from interband mixing, connecting the bound states of particles in flat Chern bands to quantum Hall effects on a torus.
Nonlinear and Nonadiabatic Transport— The nonadiabatic metric shows a profound impact on the electronic transport phenomena by modifying the wave packet velocity $\dot{x}_{i}$ in the following set of equations of motion [59]:
$\dot{q}^{i}$ $=-\frac{\partial V}{\partial x_{i}}$ (6)
@@
where the velocity $\dot{x}_{i}$ has four different contributions. The first term is the group velocity from band dispersion that was derived back to the 1930s [63, 5]. The second term is the anomalous velocity from the Berry curvature $F_{ij}=-2\text{Im}\mathcal{Q}_{ij}$ arising from the adiabatic evolution of wave packets advanced by Chang, Sundaram, and Niu [7, 8, 5]. The nonadiabatic metric leads to two velocity corrections, termed the geodesic velocity and geometric velocity separately, as detailed below.
The kinetic energy of the generalized coordinate $\bm{q}$ leads to a contribution termed the geodesic velocity, as the coefficient $\Gamma_{i,jk}$ is closely related to $\Gamma^{i}_{jk}$ that defines the geodesic motion as shown later in Eq. (10). $\Gamma_{i,jk}$ is the Christoffel symbol of the first kind:
@@
In the presence of a static electric field $\bm{\varepsilon}$, $\dot{\bm{q}}=\bm{\varepsilon}$ and $\ddot{\bm{q}}$ vanishes. The velocity thus reduces to
@@
which has been widely adopted in studying nonlinear responses [27, 64, 65, 23, 24, 25]. The velocity in Eq. (9) was originally derived by studying adiabatic evolution [27, 64], where the wave-packet wave function is expanded as an asymptotic series in the field strength. In contrast, our derivation demonstrates that the same correction can be systematically obtained by expanding the wave function as an asymptotic series in the frequency of the external fields. This perspective naturally generalizes the framework, enabling the analysis of electronic responses to a broader range of perturbations beyond static electric and magnetic fields.
@@
The second nonadiabatic velocity is new, termed the geometric velocity, as the coefficient is the nonadiabatic metric. This term is proportional to $\ddot{\bm{q}}$, the acceleration of the generalized coordinate $\bm{q}$. While $\ddot{\bm{q}}$ vanishes under static fields or is negligible in the adiabatic limit, it becomes significant in nonadiabatic responses to time-dependent, oscillating forces. For instance, in the presence of a finite-frequency electric field $\varepsilon^{i}e^{i\omega t}$, the geometric velocity acquires a nonzero value: $\ddot{q}^{i}=i\omega\varepsilon^{i}e^{i\omega t}$. At linear order in $\varepsilon^{i}$, this term gives rise to a dissipationless polarization current, which agrees with the linear response theory based on the Kubo formula [65, 66]. The full implication of the geometric velocity remains an open area for further investigation. It will be manifested in nonlinear and nonadiabatic responses to finite-frequency driving, such as below-gap light [67, 68, 69, 70] and coherent collective excitations like magnons and phonons, which are phenomena of broad relevance [71, 72, 73, 74, 75, 76].
@@
This unified framework emphasizes the fundamental role of the nonadiabatic metric $G_{ij}$. While earlier work associated $G_{ij}$ with Berry connection polarizability via expansions in field strength [27, 64, 65], our results establish its inherently nonadiabatic character, emerging from expansions in field frequency. This perspective not only provides a unified and complete understanding of nonlinear and nonadiabatic responses but also naturally generalizes to responses to non-electromagnetic drivings, such as magnonic, phononic, or thermal fields—where the nonadiabatic metric $G_{ij}$ likewise plays a pivotal role.
$\bm{q}$-Space Picture: Metric Tensor and Forced Geodesic Equation— Here we unveil the deeper geometric significance of $G_{ij}$ in momentum space, or $\bm{q}$-space, to highlight the role of $\bm{q}$ as a generalized coordinate. $G_{ij}(\bm{q})$ in the Lagrangian defines a metric tensor in the tangent vector space of the generalized velocity $\dot{\bm{q}}$ defining a line element $ds^{2}=G_{ij}dq^{i}dq^{j}$. Different from the quantum metric that defines the distance between wavefunctions in the momentum space [37], here the distance is defined for the generalized coordinate $\bm{q}$ describing the wave packet center. A $\bm{q}$-dependent $G_{ij}$ signifies a curved manifold. Conversely, in cases where $G_{ij}$ is constant across the momentum space, it corresponds to an effective mass in a flat $\bm{q}$-space, simplifying the dynamics to a Euclidean framework.
where $\Gamma^{l}_{jk}$ is the Christoffel symbol of the second kind
$G^{ji}G_{ik}=\delta^{j}_{k}$ with $\delta^{j}_{k}$ the Kronecker delta. When the right-hand side of Eq. (10) vanishes, the equation reduces to the geodesic equation for a free particle navigating a curved manifold with metric tensor $G_{ij}$. A nonzero right-hand side indicates a particle subject to forces from three sources (see Fig. 2): a scalar potential from the band dispersion $E$, a gauge field from the Berry curvature $F_{ij}$, and an external force from the wave packet velocity $\dot{x}_{j}$. This equation enables the study of gravity in condensed matter setups [77]. It is noted that Eq. (10) differs from the recent study on momentum space gravity [78, 79, 80] where $G^{ij}$ and $\ddot{\bm{q}}$ are absent and $\Gamma_{i,jk}$ instead of $\Gamma^{i}_{jk}$ is used.
When $x_{i}$ can be invertibly represented as a function of $\dot{q}^{j}$ by using $\dot{q}^{i}=-\frac{\partial V}{\partial x_{i}}$, we have
the equations of motion find their $\bm{q}$-space representation:

### v2

# Nonadiabatic Origin of Quantum-Metric Effects via Momentum-Space Metric Tensor

@@ Abstract
We reveal a fundamental geometric structure of momentum space arising from the nonadiabatic evolution of Bloch electrons. By extending semiclassical wave packet theory to incorporate nonadiabatic effects, we introduce a momentum-space metric tensor—the nonadiabatic metric. This metric gives rise to two velocity corrections, dubbed geometric and geodesic velocities, providing a unified and intuitive framework for understanding nonlinear and nonadiabatic transport phenomena beyond Berry phase effects. When the nonadiabatic metric is constant, it reduces to an effective mass, modifying flat-band electron dynamics in confining potentials. In a flat Chern band with harmonic attractive interactions, the two-body wave functions mirror the Landau-level wave functions on a torus. Furthermore, we show that the nonadiabatic metric endows momentum space with a curved geometry, recasting wave packet dynamics as forced geodesic motion.
I Introduction
@@
The adiabatic approximation, pioneered by Born and Oppenheimer, is a cornerstone of condensed matter physics. The emergence of Berry phase from adiabatic evolution [1] has profoundly shaped the study of topological phases of matter [2]. Recent interest has surged in the effects of quantum geometry in Bloch-electron dynamics beyond the Berry-phase effects, which are thought to stem from the quantum metric [3, 4, 5, 6]—the real counterpart of the Berry curvature in the quantum geometric tensor [3]. The quantum metric is connected to rich and diverse physical effects, such as nonlinear responses [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], massive particle dynamics in flat bands [21, 22, 23, 24, 25, 26, 27, 28], exciton spectra [29, 30], fractional Chern insulators [31, 32, 33, 34], Wannier function spread [35], orbital magnetic susceptibility [36, 37], among others [4, 5, 6]. Pinpointing the key physical processes in which quantum geometric effects play a significant role is essential for developing a unified understanding of their diverse manifestations.
In this work, we reveal the fundamental connection between the effects of quantum geometry and the nonadiabatic electronic dynamics. We study the nonadiabatic Bloch electron dynamics by using the semiclassical wave packet theory. The wave packet theory has connected the adiabatic Berry curvature to anomalous velocity of Bloch electrons [38, 39, 40, 41], providing a unified framework to understand Berry-phase effects in crystalline materials [2] and revolutionizing our understanding of how topological quantum materials respond to electric, magnetic, thermal, and lattice deformation stimuli [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57].
We extend the adiabatic wave packet theory to the nonadiabatic regime by including interband contributions (see Fig. 1). We begin by deriving the effective Lagrangian for the wave packet, revealing that nonadiabatic corrections emerge as a metric tensor in momentum space, which we term the nonadiabatic metric, in addition to other corrections. The relationship between this metric and the quantum metric is discussed. In the following, we elucidate on the profound implications of the nonadiabatic metric. We first discuss its impact on the wave packet velocity, which provides a unified understanding of nonlinear and nonadiabatic transport. Then we explore its effect on the quantum dynamics of Bloch electrons in flat-band systems, highlighting the interband mixing-induced correction on the bound-state spectrum.
Our results indicate that, in the dynamics of the Bloch electrons, the nonadiabatic metric is more fundamental than the quantum metric: it emerges directly from the Schrödinger equation as the leading correction beyond adiabatic evolution. It governs the dynamical behavior of Bloch electrons, providing a unified description of various dynamical phenomena. In contrast, the quantum geometric tensor is defined to describe the geometric properties of static wave functions on a parameter space, which are naturally tied to adiabatic processes [3].
@@
II Results
II.1 Formalism: Nonadiabatic Wave Packet Theory
We study the dynamics of a wave packet centered at $\bm{x}_{c}$ under slowly varying scalar potential $V(\bm{x})$. Its spatial spread is much smaller than the perturbation scale, which enables us to consider an approximated Hamiltonian local to the wave packet. The Hamiltonian can be approximated by linearizing perturbations around the wave packet center:
$\mathcal{H}\approx$ $\hat{H}_{c}+\hat{H}_{1},$ (1)
where $\hat{H}_{c}=\hat{H}(\hat{\bm{x}},\hat{\bm{p}})+V(\bm{x}_{c})$ represents the local Hamiltonian with $V(\bm{x}_{c})$ a constant and $\hat{H}$ the Hamiltonian operator of the unperturbed crystal. The corresponding Bloch bands and Bloch wave functions satisfying:
$\hat{H}|\psi_{n,\bm{q}}\rangle=E_{n}(\bm{q})|\psi_{n,\bm{q}}\rangle,$ (2)
$\hat{H}_{1}=(\hat{\bm{x}}-\bm{x}_{c})\cdot\nabla_{\bm{x}}V(\bm{x}_{c})$ is the gradient correction. Here $\nabla_{\bm{x}}V(\bm{x}_{c})$ are constants and do not induce inter-band coupling [40].
We use the following wave packet ansatz [58]:
$|\Psi(t)\rangle=\int d\bm{q}\,a(\boldsymbol{q},t)|\tilde{\psi}_{\boldsymbol{q}}\rangle,$ (3)
where $a(\boldsymbol{q},t)=|a|e^{-i\gamma}$ is a narrow distribution in momentum space centered at $\boldsymbol{q}_{c}=\int\bm{q}|a|^{2}d\bm{q}$, $|\tilde{\psi}_{\boldsymbol{q}}\rangle=e^{i\boldsymbol{q}\cdot\hat{\boldsymbol{x}}}|\tilde{u}_{\bm{q}}\rangle$ is a superposition of Bloch waves with quasimomentum $\bm{q}$, and
@@ II.1 Formalism: Nonadiabatic Wave Packet Theory
with $|u_{n}(\boldsymbol{q})\rangle$ the periodic part of the Bloch wave function $|\psi_{n,\bm{q}}\rangle$, $c_{n}$ the coefficient, and $\sum_{n}|c_{n}(\boldsymbol{q},t)|^{2}=1$. The wave packet center is at $\bm{x}_{c}=\langle\Psi|\hat{\bm{x}}|\Psi\rangle$.
@@
Here we consider an isolated band labeled by index 0. We focus on the abelian case with $|c_{0}|\approx 1$ and $c_{n}\sim 0$ for $n\neq 0$. In contrast to the adiabatic wave packet dynamics developed by Sundaram and Niu [40], which constrains the wave packet on a single energy band, here we allow for interband contributions with $c_{n}\neq 0$ (see Fig. 1). These interband components $c_{n}$ are subsequently expressed as an asymptotic series in the rate of parameter evolution such as $\dot{\bm{q}}_{c}$. This contrasts with the higher-order expansion in Ref. [8], where the wave packet is expanded as a series in the strength of external fields rather than in the rate of change of the underlying parameters.
@@
The dynamics of the wave packet is obtained by minimizing the Dirac-Frenkel action (with $\hbar=1$[59, 60]
$\mathcal{S}$ $=\int dt\langle\Psi|i\frac{d}{dt}-\mathcal{H}|\Psi\rangle=\int dtL$ (5)
$L$ $=\sum_{n}c_{n}^{*}i\dot{c}_{n}-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c}+\Lambda_{[00]}+\sum_{n\neq 0}-c_{n}^{*}c_{n}\left[\omega_{[n0]}-(\Lambda_{[nn]}-\Lambda_{[00]})\right]+(c_{0}^{*}c_{n}\Lambda_{[0n]}+c.c.)$
where $\dot{c}_{n}=dc_{n}/dt$, $\omega_{[mn]}=E_{m}-E_{n}$, $\Lambda_{[mn]}=\dot{\bm{q}}_{c}\cdot\bm{A}_{[mn]}$, and Berry connection is ${A}_{i,[mn]}=\langle u_{m}|i\partial_{{q}^{i}}|u_{n}\rangle|_{\bm{q}_{c}}$. Here $\bm{x}_{c}$, $\bm{q}_{c}$, $c_{n}$, and $c_{n}^{*}$ are dynamical variables. By writing down the equation of motion of $c_{n}$ and integrating out their dynamics [see Methods], we obtain our central result: the nonadiabatic effective Lagrangian
$\mathcal{L}_{\rm eff}$ $=-\dot{\bm{q}}_{c}\cdot\bm{x}_{c}-E_{c}(\bm{q}_{c},\bm{x}_{c})+\dot{q}_{c}^{i}{A}_{i}+\frac{1}{2}G_{ij}\dot{q}_{c}^{i}\dot{q}_{c}^{j}$ (6)
where the Einstein summation convention is used for Cartesian coordinate indices throughout the manuscript, while summation over the band indices is written explicitly. Here $E_{c}=E_{0}(\bm{q}_{c})+V(\bm{x}_{c})$ with $E_{n}$ the eigenenergy of the $n$-th band and $V(\bm{x}_{c})$ the slowly varying external potential. ${A}_{i}$ is used as shorthand for ${A}_{i}^{00}$, capturing the Berry phase effects from adiabatic evolution of Bloch states. We emphasize that this work focuses on a slowly varying scalar potential, resulting in a gradient correction $\hat{H}_{1}$ that is proportional to a constant times $\hat{\bm{x}}-\bm{x}_{c}$. This specific form of $\hat{H}_{1}$ does not induce inter-band coupling. However, in the presence of a magnetic field or inhomogeneities, such as those arising from spin textures, the nonadiabatic theory yields additional corrections to the energy $E_{c}$ and Berry connections $A_{i}$ as discussed in our follow-up work [61]. From now on, we drop the subscript $c$, using $(\bm{x},\bm{q},E)$ for $(\bm{x}_{c},\bm{q}_{c},E_{c})$.
@@
Our key contribution is the identification of a metric term beyond the Berry phase effects, characterized by the coefficients $G_{ij}$. $G_{ij}$ acts as a metric tensor in momentum space, which is termed the nonadiabatic metric, as this metric originates from nonadiabatic corrections to wave packet dynamics: $c_{n}\propto\dot{q}^{i}$ for $n\neq 0$, which reflect a finite occupation of other instantaneous eigenstates. The matrix element reads
$G_{ij}=$ $2\text{Re}\sum_{m\neq 0}\frac{A_{i,[0m]}A_{j,[m0]}}{\omega_{[m0]}}.$ (7)
@@
The metric is a gauge-invariant symmetric tensor [62]. The nonadiabatic metric is related to but different from the quantum metric $g_{ij}=\text{Re}\mathcal{Q}_{ij}$ that is the real part of the quantum geometric tensor (or the Fubini-Study metric) $\mathcal{Q}_{ij}=\sum_{m\neq n}{A_{i,[nm]}A_{j,[mn]}}$. In a two-band model with an energy gap $\Delta(\bm{q})$, $G_{ij}$ of the lower band is proportional to $g_{ij}$: $G_{ij}=\frac{2}{\Delta}g_{ij}$, which is positive-semidefinite. For the upper band, the nonadiabatic metric is negative-semidefinite. The energy denominator makes $G_{ij}$ not necessarily positive-semidefinite and thus not necessarily a Riemannian metric.
@@
Just as adiabatic wave packet theory unifies Berry phase effects through the term $\dot{q}^{i}A_{i}$ [2], the nonadiabatic corrections provide a comprehensive framework for capturing a wide array of phenomena that extend beyond the Berry phase effects. Below, we discuss the nonadiabatic effects in transport and quantum dynamics of Bloch electrons.
II.2 Nonlinear and Nonadiabatic Transport
Quantum-metric contributions have been identified in diverse electronic transport phenomena, ranging from nonlinear responses [4, 5, 6] to quantum capacitance [63]. Yet, in contrast to Berry-phase physics, where wave packet dynamics supply a unified and intuitive interpretive framework, no analogous picture exists for understanding the origin of these quantum-metric-related effects. Our nonadiabatic theory provides such a framework. Specifically, the nonadiabatic metric contributes directly to the wave-packet velocity, as expressed in the following equations [58]:
$\dot{q}^{i}$ $=-\frac{\partial E}{\partial x_{i}}$ (8)
@@ II.2 Nonlinear and Nonadiabatic Transport
where the velocity $\dot{x}_{i}$ has four different contributions. The first term is the group velocity from band dispersion that was derived back to the 1930s [64, 40]. The second term is the anomalous velocity from the Berry curvature $F_{ij}=\partial_{q^{i}}{A}_{j}-\partial_{q^{j}}{A}_{i}$ arising from the adiabatic evolution of wave packets advanced by Chang, Sundaram, and Niu [42, 43, 40]. In addition, the nonadiabatic metric leads to two new velocity corrections, termed the geodesic velocity and geometric velocity separately, as detailed below.
The nonadiabatic metric term behaves like a kinetic energy of the generalized coordinate $\bm{q}$ in a curved space. The first velocity correction from it is termed the geodesic velocity, as the coefficient $\Gamma_{i,jk}$ is closely related to $\Gamma^{i}_{jk}$ that defines the geodesic motion as shown later in equation (19). $\Gamma_{i,jk}$ is the Christoffel symbol of the first kind:
@@
In the presence of a static electric field $\bm{\varepsilon}$, $\dot{\bm{q}}=\bm{\varepsilon}$ and $\ddot{\bm{q}}$ vanishes (i.e., we set the electron charge $e$ to unity). The velocity thus reduces to
@@
which has been widely adopted in studying nonlinear responses [8, 65, 63, 4, 5, 6]. By combining the velocity with the Boltzmann distribution function $f(\bm{q})=f_{0}(\bm{q}-\tau\dot{\bm{q}})$ under the relaxation time approximation, where $f_{0}$ is the Fermi-Dirac distribution function and $\tau$ is the relaxation time, one can calculate the current $j_{i}=\int\dot{x}_{i}(\bm{q})f(\bm{q})d\bm{q}$ and thereby the linear and nonlinear conductance. It is noted that the velocity in equation (11) was originally derived by studying adiabatic evolution [8, 65], using an asymptotic expansion in external field strengths. In contrast, our derivation uses an expansion in the time derivatives of intrinsic dynamical variables, which reproduces previous results in the static field limit.
@@
The second nonadiabatic velocity is new, termed the geometric velocity, as the coefficient is the nonadiabatic metric $G_{ij}$. This term is proportional to $\ddot{\bm{q}}$, the acceleration of the generalized coordinate $\bm{q}$. While $\ddot{\bm{q}}$ vanishes under static fields or is negligible in the adiabatic limit, it becomes significant in nonadiabatic responses to time-dependent, oscillating forces. For instance, in the presence of a finite-frequency electric field $\varepsilon^{i}=\varepsilon_{0}^{i}e^{i\omega t}$ with $\varepsilon_{0}^{i}$ the amplitude and $\omega$ the frequency, we find that $\dot{q}^{i}={\varepsilon}^{i}$ and thus the geometric velocity acquires a nonzero value: $G_{ij}\ddot{q}^{i}=G_{ij}\dot{\varepsilon}^{i}$. At linear order in $\varepsilon^{i}_{0}$, this term gives rise to a dissipationless polarization current $j_{i}\sim\int G_{ij}f(\bm{q})d\bm{q}\dot{\varepsilon}^{j}$, which agrees with the linear response theory based on the Kubo formula (see Sec. III in the Supplemental Material [58][63, 66]. Specifically, in an insulator, the integral corresponds to the electric susceptibility $\chi_{ij}=\int G_{ij}d\bm{q}$ [63]. This is in contrast to the quantum metric, the integral of which in the Brillouin zone is related to the spread of the Wannier function [35]. The full implication of the geometric velocity remains an open area for further investigation. It will be manifested in nonlinear and nonadiabatic responses to finite-frequency driving, such as below-gap light [67, 68, 69, 70], second-harmonic generation [71], and coherent collective excitations like magnons and phonons, which are phenomena of broad relevance [72, 73, 74, 75, 76, 77].
@@
This unified framework emphasizes the fundamental role of the nonadiabatic metric $G_{ij}$. While earlier work associated $G_{ij}$ with Berry connection polarizability via expansions in strengths of external fields [8, 65, 63], our results establish its inherently nonadiabatic nature, emerging from expansions in the changing rates of intrinsic dynamical parameters. This perspective not only offers a unified understanding of nonlinear and nonadiabatic responses, but also naturally extends to responses to non-electromagnetic fields, such as lattice dynamics and spin precessions, where the nonadiabatic metric $G_{ij}$ likewise plays a pivotal role, a feature not captured by previous theories.
II.3 Effective Mass Corrections for Bound States
The re-quantization of semiclassical wave packet dynamics is fruitful in describing various quantum phenomena such as Landau levels of Bloch electrons [41] and Berry curvature effects on exciton spectrum [30] and Wigner crystal phases [78]. While the importance of adiabatic Berry phase on those phenomena has attracted much attention in emerging material systems, such as transition-metal dichalcogenides, gapped graphenes, Landau levels, and moiré systems [30, 29, 79, 80, 78, 81, 82, 83], the effects of nonadiabatic corrections are yet to be explored.
We focus on the lowest Landau level as a concrete example. The simplicity enables analytical insights that can be generalized qualitatively to other flat-band systems. For the lowest Landau level, the Berry curvature $F_{ij}$ and nonadiabatic metric $G_{ij}$ are constants. Consider a rectangular Brillouin zone with dimensions $L_{q_{x},q_{y}}=2\pi/L_{x,y}$ where $L_{q_{x},q_{y}}$ defines the $\bm{q}$}-space size while $L_{x,y}$ are the dimensions of real-space unit cell with one magnetic flux quantum. The Berry curvature is uniform and its integral over the Brillouin zone corresponds to the Chern number $C$, i.e., $F=\frac{2\pi C}{L_{q_{x}}L_{q_{y}}}$. The nonadiabatic metric reads $G_{ij}=G_{0}\mathcal{I}_{ij}$ with $\mathcal{I}$ the identity matrix and $G_{0}=\frac{|F|}{\Delta}$ where $\Delta$ is the energy spacing between the lowest and the first Landau levels (see Supplemental Material) [58].
We consider the wave packet dynamics in a harmonic potential that can arise from external electrostatic fields or Coulomb interactions in Wigner crystals [78, 81]
$V(\bm{x})=\frac{\kappa}{2}\bm{x}^{2}.$ (12)
Thus, we have the equation of motion $\dot{q}^{i}=-\kappa x^{i}$ and thereby $\dot{x}^{i}=-\frac{1}{\kappa}\ddot{q}^{i}$. Here $x^{i}=\delta^{ij}x_{j}$ and $\delta^{ij}$ the Kronecker delta. The equation of $\bm{q}$ can be expressed as a forced geodesic equation in the momentum space ($\bm{q}$-space):